OpenStudy (sidsiddhartha):

ok need a little help : need to solve these--\[\large \frac{ d^2x }{ dt^2 }-y=2x+2t\\\large \frac{ d^2x }{ dt^2 }+4\frac{ dy }{ dt }=3y\]

OpenStudy (sidsiddhartha):

i was trying to solve it using CF and PI

OpenStudy (perl):

do you have notes on this

OpenStudy (sidsiddhartha):

my work: \[y=\frac{ d^2x }{ dt^2 }-2x-2t\] now putting this one in the second equation -- \[\frac{ d^2x }{ dt^2 }+4[\frac{ d^3x }{ dt^3 }-2\frac{ dx }{ dt }-2]=3[\frac{ d^2x }{ dt^2 }-2x-2t]\\2\frac{ d^3x }{ dt^3 }-\frac{ d^2x }{ dt^2 }-4\frac{ dx }{ dt }+3x=4-3t\] so the auxiliary eq. will be \[(2D^3-D^2-4D+3)x=4-3t\\2m^3-m^2-4m+3=0\]

OpenStudy (sidsiddhartha):

solving it \[(m-1)^2(2m+3)=0\\m=1,1,-3/2\\so~~C.F=(c_1+c_2t)e^t+c_3e^{\frac{ -3 }{ 2 }t}\] am i right so far?

OpenStudy (sidsiddhartha):

hartnn?

hartnn (hartnn):

looks good, bit rusty on solving D.Es, recently only using laplace transform to solve DEs

OpenStudy (sidsiddhartha):

yeah LT is easier way, but i have to use these now im going ahead with the PI

hartnn (hartnn):

we can always verify the answer later using LT

OpenStudy (sidsiddhartha):

now in the proper integral part--\[P.I=\frac{ 1 }{ 2D^3-D^2-4D +3}(4-3t)\] right?

OpenStudy (sidsiddhartha):

now what to do in the next step ,i dont see any way to use some formulaes :(

hartnn (hartnn):

use partial fractions ? 1/(D-1)^2 (2D+3) = A/(D-1) + B/(D-1)^2 + C/(2D+3)

hartnn (hartnn):

then apply each of them individually to 4-3t

OpenStudy (sidsiddhartha):

yes good idea i can use partial or binomial ,ok wait a min

OpenStudy (sidsiddhartha):

\[PI=\frac{ 1 }{ 3 }[1+\frac{ 2D^3-D^2-4D }{ 3}]^{-1}(4-3t)\\=\frac{ 1 }{ 3 }[1-\frac{ 2 }{ 3 }D^3+\frac{ D^2 }{ 3 }+\frac{ 4 }{ 3}D+....](4-3t)\] ok?

hartnn (hartnn):

right, yes

hartnn (hartnn):

you can drop D^2 and D^3 terms...

OpenStudy (sidsiddhartha):

yeah it gives me \[PI=1/3(4-3t-4)=-t\]

OpenStudy (sidsiddhartha):

yeah \[so\\x=(c_1+c_2t)e^t+c_3e^{\frac{ -3 }{ 2 }t}-t\] thats the answer thanks :)

hartnn (hartnn):

welcome ^_^

OpenStudy (sidsiddhartha):

lets try LT

OpenStudy (sidsiddhartha):

\[first~eq\\s^2X(s)-Y(s)=2X(s)+2/s^2............(1)\\second~one\\s^2X(s)+4Y(s)=3Y(s)...............(2)\] uhh i think it will more complicated

OpenStudy (sidsiddhartha):

*mistake \[second ~one\\s^2X(s)+4sY(s)=3Y(s)\]

hartnn (hartnn):

right, more complicated than the binomial method u used, less if you had used partial fractions as i suggested...

OpenStudy (anonymous):

@sidsiddhartha is this a system of ODEs, or two separate ones?

OpenStudy (sidsiddhartha):

yes system of ODEs @SithsAndGiggles

OpenStudy (anonymous):

Ok, the reason I ask is because you're only given one second order derivative. I'm not familiar with the methods, or perhaps their names (PI and CF?). It looks like one of them might be some form of elimination where you substitute one eq. into the other. Another approach would be cumbersome, but it should work. It would involve order-reduction of the given system. Let \[\begin{pmatrix}x\\y\\x'\\y'\end{pmatrix}=\begin{pmatrix}u_1\\u_2\\u_3\\u_4\end{pmatrix}\] Then you have \[\begin{cases} x''=2x+y+2t\\ x''+4y'=3y&\text{(are there really two of }x''?) \end{cases}\] changed to \[\begin{cases} u_1'=u_3\\ u_2'=u_4\\ u_3'=2u_1+u_2+2t\overbrace{=}^{?}3u_2-4u_4\\ u_4'=??? \end{cases}\] As you can see, though, I'm not sure how to deal with there being two \(x''\)s...

OpenStudy (sidsiddhartha):

yeah its like separation of variables and yes that is a problem thats why i used CF and Pi method u can check this out

OpenStudy (anonymous):

That's a nice set of notes. You hardly see any that give exercises WITH solutions. Thanks for the upload!

OpenStudy (sidsiddhartha):

ur welcome :)